OpenStudy (anonymous):

Four polynomial functions are shown below: f(x) = x3 + 3x2 - 4x - 12 g(x) = x3 - 3x - 2 p(x) = x2 + 3x + 2 t(x) = x2 + x - 2 Which of the following functions has -1 and -2 as its zeros? f(x) g(x) p(x) t(x)

OpenStudy (solomonzelman):

you can plug in -1 and -2 into each of the function too see in which function will both x=-1 and x=-2 result in f(x)=0. OR, you can solve each one. What do you prefer?

OpenStudy (anonymous):

well for starters since it's 2 zeros then it has to be p(x) or t(x) so p(x)= x²+3x+2 = (x+1) (x+2) t(x)= x²+x-2 = (x-1) (x+2) so the answer is p(x) because x+1= 0 x=-1 and x+2 = 0 x=-2

OpenStudy (anonymous):

p(x)